### v1

# LRQ-DiT: Log-Rotation Post-Training Quantization of Diffusion Transformers for Text-to-Image Generation

Abstract
Diffusion Transformers (DiTs) have achieved impressive performance in text-to-image generation. However, their high computational cost and large parameter sizes pose significant challenges for usage in resource-constrained scenarios. Post-training quantization (PTQ) is a promising solution to reduce memory usage and accelerate inference, but existing PTQ methods suffer from severe performance degradation under extreme low-bit settings. We identify two key obstacles to low-bit post-training quantization for DiT models: (1) model weights follow a Gaussian-like distribution with long tails, causing uniform quantization to poorly allocate intervals and leading to significant errors; (2) two types of activation outliers: (i) Mild Outliers with slightly elevated values, and (ii) Salient Outliers with large magnitudes concentrated in specific channels, which disrupt activation quantization. To address these issues, we propose LRQ-DiT, an efficient and accurate PTQ framework. We introduce Twin-Log Quantization (TLQ), a log-based method that aligns well with the weight distribution and reduces quantization errors. We also propose an Adaptive Rotation Scheme (ARS) that dynamically applies Hadamard or outlier-aware rotations based on activation fluctuation, effectively mitigating the impact of both types of outliers. We evaluate LRQ-DiT on PixArt and FLUX under various bit-width settings, and validate the performance on COCO, MJHQ, and sDCI datasets. LRQ-DiT achieves low-bit quantization of DiT models while preserving image quality, outperforming existing PTQ baselines.
1 Introduction
Figure 1: The overall framework of LRQ-DiT. (a) LRQ-DiT proposes the twin-log quantization (TLQ) method to address imbalanced weight distribution and excessive quantization errors. Compared to uniform quantization, TLQ significantly enhances the practicality of low-bit quantization. (b) LRQ-DiT utilizes the metric J to identify mild and salient outliers in activations, and proposes an adaptive rotation scheme that applies customized rotations to suppress the impact of each type of outlier.
Diffusion models have achieved remarkable success in text-to-image generation, with recent advances expanding their applications across vision domains (Esser et al. 2024; Podell et al. 2023; Xie et al. 2024; Lin et al. 2025; Jiang et al. ; Zhou et al. 2024). Notably, DiTs replace the traditional U-Net backbone with the Transformer architecture, significantly improving scalability and computational efficiency. DiTs are becoming dominant in tasks such as image generation (Hatamizadeh et al. 2024; Chen et al. 2024b), video generation (Zhang et al. 2025), and image super-resolution. Despite their performance, the large number of parameters in DiT models poses serious challenges for deployment on resource-constrained edge devices, especially in terms of memory consumption and inference latency.
Model compression (Liu et al. 2018; Hinton, Vinyals, and Dean 2015; Lin et al. 2024a) offers a promising acceleration solution, with quantization being one of the most effective techniques. By converting high-precision floating-point values into low-bit integers, quantization reduces both storage and computation costs, and has been widely adopted in large-scale Transformers (Yuan et al. 2022; Li et al. 2023b). PTQ has become increasingly popular for large models, such as LLMs and diffusion transformers, due to its training-free nature (Yao et al. 2022; Shao et al. 2023). Existing PTQ methods for DiTs typically focus on moderate configurations such as W4A8 and W8A8, but performance under lower-bit settings remains limited. PTQ4DiT (Wu et al. 2024a) and Q-DiT (Chen et al. 2025) incorporate time-step-specific activation patterns to achieve W4A8 quantization on DiT-XL/2, yet they exhibit performance degradation when applied to more advanced DiT variants such as PixArt and FLUX. ViDiT-Q (Zhao et al. 2024) proposes a dynamic quantization strategy, and SVDQuant (Li et al. 2024b) addresses outliers via kernel-level protection, both achieving strong performance at the 4 bit weight quantization. However, achieving accurate quantization for DiTs remains a significant challenge, particularly due to the difficulty in preserving precision for both weights and activations under extreme low-bit constraints.
Our preliminary experiments identify two key challenges for low-bit quantization of DiT models: (1) Uniform quantizers fundamentally mismatch the Gaussian-like weight distribution of DiTs. We observe that almost all weight matrices in DiT models follow a Gaussian-like distribution with long tails on both sides. The Uniform quantization fails to allocate sufficient quantization intervals to the dense central region and ignores the long-tail extremes, resulting in large quantization errors under low-bit settings. (2) DiT activations contain two distinct types of outliers that disrupt low-bit quantization. The majority of outliers in activations are Mild Outliers, with values slightly exceeding normal activation levels and rarely surpassing 5. A smaller portion is Salient Outliers, represented as high-magnitude activations concentrated in specific channels. These outliers pose significant challenges for low-bit activation quantization.
To address the challenges identified above, we propose LRQ-DiT, which integrates two core components: a twin-log quantization (TLQ) method for Gaussian-like model weights, and an adaptive rotation scheme (ARS) for accurate activations quantization. First, to better match the weight distribution, we introduce twin-log quantization. We apply a logarithmic transformation to model weights and quantize the positive and negative parts separately. A simple search-based clipping strategy is further employed to suppress extreme values in the long-tail regions. Second, we propose an adaptive rotation scheme to handle Mild and Salient activation outliers. To balance the accuracy and efficiency, we define an adaptive metric to measure the fluctuation level of activations across layers. Layers with low fluctuation are processed using Hadamard rotation, while layers with high fluctuation adopt dual transformations, combining greedy outlier-aware rotation with channel-wise permutation for enhanced precision. This design enables LRQ-DiT to effectively mitigate both types of outliers, while maintaining high precision and deployment efficiency.
Our main contributions are summarized as follows:
  • We observe that uniform quantization struggles to adapt to the Gaussian-like distribution of weights, leading to significant quantization errors. To address this issue, we propose twin-log quantization to resolve this issue.
  • We leverage the metric to identify both mild and salient outliers in activations and propose an adaptive rotation scheme to perform targeted rotations for these two types of outliers, achieving effective outlier suppression.
  • Extensive quantization experiments conducted on PixArt and FLUX models under varying bit settings and datasets demonstrate the effectiveness of our method LRQ-DiT.
2 Related Works
2.1 Text-to-Image Models
Text-to-image generation has made remarkable progress, largely driven by advancements in model architecture. Stable Diffusion (Rombach et al. 2022) bridges the semantic gap between text and image modalities, and its successor, Stable Diffusion 3 (Esser et al. 2024), further enhances multi-modal fusion. Subsequent models such as SDXL (Podell et al. 2023) and its variants provide a strong foundation for efficient and high-quality text-to-image generation. More recently, transformer-based architectures have replaced U-Net in diffusion models, opening new possibilities for downstream applications. PixArt-$\alpha$ (Chen et al. 2023) reduces training cost while maintaining image quality, and PixArt-$\Sigma$ (Chen et al. 2024a) significantly improves 4K image generation efficiency. FLUX (Labs 2024) (primarily FLUX.1-schnell and FLUX.1-dev) further demonstrates the capability of DiT-based models to generate high-resolution, high-fidelity images. In this work, we focus on diffusion transformers and propose LRQ-DiT, a low-bit weight–activation quantization framework tailored for DiT models to enhance the quality of text-to-image generation.
2.2 Model Quantization
Quantization reduces model memory by converting high-precision floating-point weights to low-bit integers, while activation quantization accelerates inference. Post-training quantization (PTQ) (Yang et al. 2024b, a) is also widely used in large-scale pre-trained models as it avoids costly retraining. For large language models (LLMs), outliers (Dettmers et al. 2022; Sun et al. 2024), the values significantly larger than most activations, pose a major challenge. SmoothQuant (Xiao et al. 2023) shifts quantization difficulty to weights for better 8-bit performance, while rotation-based methods (Ashkboos et al. 2024; Liu et al. 2024b; Lin et al. 2024b) balance outliers via rotation transformations, enabling competitive 4-bit weight–activation quantization. Applying these techniques to diffusion models is nontrivial due to time-step-dependent features. For diffusion models (DMs), Q-Diffusion (Li et al. 2023a) and PTQ4DM (Shang et al. 2023) considered the influence of time-steps when determining quantization parameters, achieving effective 8-bit quantization, which also stimulated subsequent works (He et al. 2023; Huang et al. 2024; Wang et al. 2024). However, there are still differences in the structure between DiTs and DMs. Recent works address these variations by assigning channel-wise quantization parameters (Chen et al. 2025), reallocating salient channels across time steps (Wu et al. 2024a), protecting activation outliers with low-rank branches (Li et al. 2024b), or using fine-grained grouping with dynamic quantization (Zhao et al. 2024). Preliminary progress has also been made in quantizing video generation models (Feng et al. 2025; Li et al. 2025). However, these methods still degrade significantly under extreme low-bit settings, motivating us to explore the underlying challenges and propose LRQ-DiT.
3 Method
3.1 Preliminaries
Quantization.
The uniform quantizer maps floating-point data $\bm{x}$ to integer data $\bm{x}_{q}$, with a given bitwidth $b$. The quantization and dequantization process can be expressed as:
$\bm{x}_{q}$ $=\text{clamp}\left(\left\lfloor\frac{\bm{x}}{s}\right\rceil-z,0,2^{b}-1\right),$ (1)
${\bm{x}_{f}}$ $=s\cdot(\bm{x}_{q}+z)\approx\bm{x},$
$\text{where}~{s}$ $=\frac{\bm{x}_{max}-\bm{x}_{min}}{2^{b}-1},{z}=\left\lfloor\frac{\bm{x}_{min}}{s}\right\rceil.$
Among them, $\left\lfloor\cdot\right\rceil$ is the rounding function, clamp()($\cdot$) is the function that limits the value to the range of $0,2^{b}-1]$, $s$ denotes the scaling factor, and $z$ represents the zero-point.
Rotation Transformation.
Recently, some rotation-based methods (Ashkboos et al. 2024; Liu et al. 2024b; Lin et al. 2024b) eliminate outliers in LLM activations via the rotation matrix $\mathbf{R}$ along with corresponding weight transformations.
$\mathbf{Y}$ $=\mathbf{X}\mathbf{W}^{\top}=(\mathbf{X}\mathbf{R})(\mathbf{R}^{\top}\mathbf{W}^{\top}),$ (2)
where $~\mathbf{R}\mathbf{R}^{\top}=I\quad\text{and}\quad|\mathbf{R}|=1.$
Figure 2: Comparison of 3-bit quantization errors of PixArt weights ($W_{q}-W||_{2}$) under different quantization methods.
3.2 Twin-Log Quantization for Weights
OB 1. Uniform quantization mismatches the Gaussian distribution of weights, leading to significant quantization errors and limiting the performance of low-bit quantized models.
Motivation.
Inspired by previous studies (Li et al. 2025; Feng et al. 2025) and our experiments, we visualize the weight distributions of DiTs and observe that most weight matrices follow an approximately zero-mean Gaussian-like distribution with long tails on both sides, as shown in Figure 1(a). Existing works (Zhao et al. 2024; Wu et al. 2024a) primarily adopt uniform quantizers for weight quantization. However, our preliminary experiments reveal that 3-bit uniform quantization introduces a large error ($W_{q}-W||_{2}$) on PixArt weights, illustrated in Figure 2. This explains the severe accuracy degradation under low-bit settings. We attribute this to two factors: ➊ uniform quantizer cannot allocate sufficient intervals to the dense central region, limiting the expressive power of the majority of parameters; and ➋ uniform quantizer fails to handle the long-tail regions, where parameters are few but have large magnitudes.
Twin-Log Quantization (TLQ)
. To address the limitations of uniform quantization, we propose a twin-log quantization (TLQ) method designed to better fit the Gaussian-like weight distribution. Regarding issue ➊, we introduce a log-based quantizer to allocate more quantization intervals to the dense central region. Specifically, we first apply a logarithmic transformation to obtain $\bm{W}^{\prime}=\log_{2}(|\bm{W}|)$. The transformed weight matrix $\bm{W}^{\prime}$ is then divided into positive and negative parts using position masks $\bm{M^{+}}$ and $\bm{M^{-}}$:
$\bm{W}^{+}=\bm{W}^{\prime}\cdot\bm{M^{+}},\quad\bm{W}^{-}=\bm{W}^{\prime}\cdot\bm{M^{-}}.$ (3)
where $\bm{M^{+}}$ and $\bm{M^{-}}$ indicate the positions of positive and negative values in $\bm{W}$, respectively. We quantize the positive and negative parts separately, and the $b$-bit quantization is finally performed as:
$\bm{W}_{q}^{+}$ $=\text{clamp}\left(\left\lfloor\frac{\bm{W}^{+}}{s^{+}}\right\rceil-z^{+},0,2^{b-1}-1\right),$ (4)
$\bm{W}_{q}^{-}$ $=\text{clamp}\left(\left\lfloor\frac{\bm{W}^{-}}{s^{-}}\right\rceil-z^{-},0,2^{b-1}\right),$
$\bm{W}_{q}^{~}$ $=\bm{W}_{q}^{+}-\bm{W}_{q}^{-}.$
where $s^{+}$ and $s^{-}$ denote the scaling factors, and $z^{+}$ and $z^{-}$ are zero-points for the positive and negative parts, respectively. This log-based design also naturally de-emphasizes the long-tail regions of the weight distribution, where parameters are few but have large magnitudes. As shown in Figure 1(a), this simultaneously addresses issue ➋.
Search-based Clipping.
Inspired by learnable weight clipping (LWC) (Shao et al. 2023), we adopt a search-based clipping strategy to further suppress quantization errors in long-tail regions. Specifically, two hyperparameters, $\alpha$ and $\beta$, are introduced to clip the extreme values on both sides of the weight distribution, and their optimal values are determined via a simple grid search. The clipping-based scaling factors and zero-points are calculated as:
$s^{+}$ $=\frac{\alpha\cdot\bm{W}_{\max}^{+}-\bm{W}_{\min}^{+}}{2^{b-1}-1},\quad z^{+}=\left\lfloor\frac{\bm{W}_{\min}^{+}}{s^{+}}\right\rceil,$ (5)
$s^{-}$ $=\frac{\beta\cdot\bm{W}_{\max}^{-}-\bm{W}_{\min}^{-}}{2^{b-1}},\quad z^{-}=\left\lfloor\frac{\bm{W}_{\min}^{-}}{s^{-}}\right\rceil.$
The de-quantized weights are reconstructed as:
$\bm{W}_{f}=2^{\,s^{+}(\bm{W}_{q}^{+}+z^{+})}\cdot\bm{M}^{+}-2^{\,s^{-}(\bm{W}_{q}^{-}+z^{-})}\cdot\bm{M}^{-}.$ (6)
This clipping strategy effectively suppresses extreme long-tail weights while preserving the dense regions of the distribution, further improving low-bit quantization performance.
Hardward Implementation.
Since the exponential operations in Equation (6) are not integer-friendly, existing TLQ still has potential for further acceleration in practical deployment. To address this, we design a hardware-oriented acceleration pipeline by decomposing the exponential term into an integer part and a residual part. Specifically, for the quantized weights, we split the exponents into:
$f^{+}$ $=\left\lfloor s^{+}(\bm{W}_{q}^{+}+z^{+})\right\rfloor,\quad r^{+}=s^{+}(\bm{W}_{q}^{+}+z^{+})-f^{+},$ (7)
$f^{-}$ $=\left\lfloor s^{-}(\bm{W}_{q}^{-}+z^{-})\right\rfloor,\quad r^{-}=s^{-}(\bm{W}_{q}^{-}+z^{-})-f^{-},$
$\bm{W}_{f}$ $=2^{f^{+}}\cdot 2^{r^{+}}\cdot\bm{M}^{+}-2^{f^{-}}\cdot 2^{r^{-}}\cdot\bm{M}^{-}.$
We then introduce an integerization factor $2^{-I}$ (e.g., 1/64, 1/128, 1/256) (Wu et al. 2024b) to approximate the residual exponents by integers while minimizing quantization error:
$2^{r^{+}}$ $\approx\mathbb{I}^{r+}\cdot 2^{-I}=\left\lfloor\frac{2^{r^{+}}}{2^{-I}}\right\rceil\cdot 2^{-I},$ (8)
$2^{r^{-}}$ $\approx\mathbb{I}^{r-}\cdot 2^{-I}=\left\lfloor\frac{2^{r^{-}}}{2^{-I}}\right\rceil\cdot 2^{-I}.$
Combining Equations (7) and (8), the weight–activation multiplication can be expressed as:
$\bm{W}\bm{X}$ $=~\left(2^{f^{+}}\mathbb{I}^{r+}\bm{M}^{+}-2^{f^{-}}\mathbb{I}^{r-}\bm{M}^{-}\right)2^{-I}\mathbb{I}^{X}s^{X}$ (9)
$=~~~[(\mathbb{I}^{r+}\mathbb{I}^{X})<<(f^{+}-I)]\bm{M}^{+}s^{X}$
$\quad-[(\mathbb{I}^{r-}\mathbb{I}^{X})<<(f^{-}-I)]\bm{M}^{-}s^{X}.$
$\mathbb{I}^{X}$ and $s^{X}$ represent the integers and scales of activations. This formulation enables efficient hardware execution: the SHIFT operation accelerates power-of-two multiplications, and mask multiplications are implemented via fast bitwise AND operations. Therefore, TLQ can be deployed with lower latency and reduced computational cost.
3.3 Adaptive Rotation Scheme for Activations
OB 2. We identify two types of activation outliers in DiTs: Mild Outliers with slightly elevated values and Salient Outliers with larger magnitudes in specific channels.
Motivation.
We observe two distinct types of activation outliers in DiT models. Here, outliers refer to values significantly larger than typical activations. Based on their magnitude, we categorize them as  ➊ Mild Outliers and  ➋ Salient Outliers. As shown in Figure 1(b), most outliers slightly exceed typical activation values and rarely surpass 5, which we term Mild Outliers. In contrast, Salient Outliers appear in specific channels with much larger magnitudes. For instance, in the blocks.26.attn2.to_\_q layer of PixArt, there is a notable phenomenon of numerical outliers: 3.6% of the activation values exceed 10, while a small but critical subset (0.1%) surpasses 100, with the peak value reaching as high as 245. These disproportionately large activations pose significant challenges for low-bit weight-activation quantization. Besides, these Salient Outliers are prompt-independent and consistently observed across models such as PixArt and FLUX, with more pronounced patterns in PixArt. Notably, our observation differs from the Massive Outliers reported in LLMs (Sun et al. 2024; Liu et al. 2024a), which typically appear in a few specific tokens. In contrast, both Mild and Salient Outliers in DiT models are distributed across all tokens but confined to specific channels.
Model Precision Method COCO dataset MJHQ dataset sDCI dataset
FID $\downarrow$ IR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ FID $\downarrow$ IR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ FID $\downarrow$ IR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$
W3A4 Smooth* 437.47 -2.28 0.08  6.92 444.72 -2.28 0.09  6.74 429.29 -2.28 0.09  6.97
QuaRot* 274.75 -2.03 0.36 10.28 223.76 -1.72 0.40 11.14 261.14 -1.94 0.31 10.71
ViDiT-Q* 210.55 -1.33 0.35 10.77 200.83 -1.53 0.36 11.01 239.69 -1.66 0.30 10.62
Ours  90.15  0.43 0.52 13.47  72.29  0.86 0.52 14.02  89.06  0.47 0.43 13.46
W3A6 Smooth* 496.73 -2.28 0.09  7.31 539.20 -2.28 0.09  7.42 505.05 -2.28 0.11  7.29
QuaRot* 152.06 -1.22 0.46 10.87 107.67 -0.39 0.49 11.73 134.60 -1.09 0.40 11.30
ViDiT-Q* 115.63 -0.30 0.42 11.41  99.05 -0.22 0.46 11.86 124.75 -0.65 0.37 11.54
Ours  52.97  0.79 0.53 13.47  47.37  1.10 0.54 14.16  53.61  0.68 0.45 13.24
W3A8 Smooth* 501.06 -2.28 0.09  7.25 569.98 -2.28 0.09  7.35 510.94 -2.28 0.11  7.31
QuaRot* 144.83 -1.15 0.47 10.89 103.68 -0.32 0.49 11.75 129.51 -1.04 0.40 11.33
ViDiT-Q* 111.21 -0.25 0.43 11.44  96.71 -0.17 0.46 11.88 120.07 -0.59 0.38 11.65
PixArt-$\Sigma$ Ours  51.06  0.82 0.52 13.36  47.56  1.12 0.58 14.51  53.56  0.80 0.45 13.51
20 Steps W4A4 Smooth* 392.64 -2.28 0.16  7.01 397.50 -2.28 0.13  6.25 389.65 -2.28 0.15  6.80
QuaRot*  86.03  0.65 0.47 13.01  67.24  1.02 0.51 14.17  81.64  0.62 0.40 13.05
ViDiT-Q*  67.65  0.78 0.53 14.40  60.08  1.12 0.35 10.97  68.61  0.81 0.25 10.05
Ours  65.63  0.83 0.54 14.47  53.87  1.16 0.55 15.44  61.94  0.82 0.43 13.77
W4A6 Smooth*  97.31  0.20 0.35 11.24  82.86  0.42 0.35 11.26  94.96 -0.09 0.32 10.90
QuaRot*  42.96  0.86 0.56 13.91  38.97  1.21 0.60 15.05  43.02  0.93 0.49 13.95
ViDiT-Q*  38.08  0.90 0.61 15.14  34.99  1.25 0.64 16.21  38.12  0.97 0.55 15.05
Ours  35.83  0.95 0.63 15.56  32.79  1.27 0.66 16.66  35.57  1.00 0.57 15.69
W4A8 Smooth*  72.51  0.48 0.36 11.33  67.76  0.68 0.36 11.32  72.14  0.23 0.33 11.06
QuaRot*  41.78  0.83 0.56 13.94  37.90  1.21 0.61 15.08  41.80  0.93 0.50 14.01
ViDiT-Q*  36.96  0.88 0.62 15.17  34.23  1.24 0.65 16.25  37.41  0.98 0.56 15.13
Ours  35.16  0.95 0.64 15.64  32.39  1.27 0.66 16.53  34.14  0.99 0.59 15.82
Table 1: Performance of LRQ-DiT (PixArt-$\Sigma$) text-to-image generation on different datasets. FID is calculated based on the images generated by FP16 inference. To ensure fairness in the comparison, we align the experimental setup and rerun SmoothQuant, QuaRot, ViDiT-Q, denoted as Method*. We follow the setting of ViDiT-Q to preserve FP precision for certain sensitive and parameter small layers (e.g., embed, norm_\_out, proj_\_out, adaln_\_single, caption_\_projection), denoted as [WnAn].
Adaptive Rotation Scheme (ARS)
. Recently, rotation-based transformations have been proven effective in mitigating outliers in LLMs. The use of Hadamard matrices $\mathbf{R}_{H}\in\{+1,-1\}^{n\times n}$ provides a simple and efficient way to smooth outliers (Ashkboos et al. 2024; Chee et al. 2023; Tseng et al. 2024). Moreover, DuQuant (Lin et al. 2024b) further enhances performance on massive outliers by constructing rotation matrices based on the largest outlier channel and introducing permutation transformations to balance the activation distribution. The Hadamard rotation is lightweight and straightforward, while DuQuant’s operations are more effective for handling high-magnitude outliers.
For DiT models, where Mild and Salient Outliers appear in different linear layers, we propose an adaptive rotation scheme tailored to different activation inputs. Specifically, we introduce an adaptive metric $\mathbf{J}$ to measure the fluctuation degree in layer activations with shape $BNC$:
$\mathbf{J}$ $=\frac{\|\mathbf{X}\|_{F}}{\sqrt{BNC}}$ (10)
Here, $\|\mathbf{X}\|_{F}$ captures the overall activations’ magnitude, and the term $\sqrt{BNC}$} normalizes the scale across layers. Based on this metric, we apply Hadamard rotation to activation matrices with $\mathbf{J}$ lower than a predefined threshold, while activations with larger $\mathbf{J}$ adopt DuQuant’s dual transformations.
[…]
Model Precision Method COCO dataset MJHQ dataset sDCI dataset
FID $\downarrow$ IR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ FID $\downarrow$ IR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ FID $\downarrow$ IR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$
W3A4 Smooth* 321.59 -2.26 0.40  7.28 350.92 -2.27 0.35  4.53 329.29 -2.26 0.39  8.22
QuaRot* 292.01 -2.25 0.36 10.93 280.91 -2.19 0.36 11.89 271.63 -2.20 0.35 11.27
ViDiT-Q* 245.37 -2.04 0.35 10.37 232.04 -1.86 0.31 11.89 216.56 -2.03 0.32 11.61
Ours  90.36  0.01 0.49 12.82  75.51  0.48 0.54 13.90  87.06 -0.25 0.48 13.22
W3A6 Smooth* 299.82 -2.26 0.40  8.15 293.05 -2.26 0.41  7.07 327.55 -2.27 0.39  7.28
QuaRot* 186.82 -2.08 0.43 11.10 176.33 -1.90 0.46 12.10 170.67 -2.03 0.42 11.47
ViDiT-Q* 151.59 -1.46 0.43 10.78 124.88 -0.87 0.43 12.45 138.73 -1.63 0.39 11.85
Ours  57.76  0.31 0.53 12.86  51.95  0.85 0.56 14.12  61.41  0.03 0.49 13.36
W3A8 Smooth* 307.17 -2.26 0.39  7.97 303.78 -2.26 0.41  6.61 316.17 -2.27 0.39  7.34
QuaRot* 183.38 -2.06 0.43 11.09 170.95 -1.89 0.47 12.10 167.23 -2.03 0.42 11.47
ViDiT-Q* 149.14 -1.43 0.41 10.43 121.65 -0.82 0.44 12.47 136.03 -1.61 0.39 11.84
PixArt-$\alpha$ Ours  56.95  0.30 0.52 12.83  50.84  0.87 0.57 14.15  61.17  0.05 0.49 13.35
20 Steps W4A4 Smooth* 375.94 -2.25 0.42  9.85 367.77 -2.26 0.43 10.09 385.19 -2.26 0.39  9.60
QuaRot* 120.81  0.26 0.19 9.86  79.17  0.79 0.33 12.87 135.87 -0.13 0.25 11.54
ViDiT-Q*  72.12  0.64 0.49 13.78  60.45  1.12 0.47 14.59  78.69  0.51 0.41 13.72
Ours  61.83  0.77 0.51 14.30  50.18  1.24 0.56 15.41  74.33  0.55 0.41 13.85
W4A6 Smooth* 405.53 -2.18 0.13  8.33 336.22 -1.97 0.13  9.08 324.68 -1.94 0.14  8.87
QuaRot*  42.71  0.75 0.58 14.63  38.04  1.27 0.61 15.62  45.01  0.70 0.54 14.68
ViDiT-Q*  40.92  0.74 0.60 14.62  37.96  1.28 0.62 15.84  44.27  0.73 0.55 14.87
Ours  36.58  0.75 0.62 15.47  33.92  1.32 0.65 16.50  40.74  0.77 0.56 15.50
W4A8 Smooth* 407.81 -2.18 0.12  8.32 354.63 -2.02 0.13  8.93 334.11 -1.96 0.13  8.71
QuaRot*  41.41  0.77 0.59 14.67  37.48  1.27 0.62 15.67  44.11  0.68 0.55 14.73
ViDiT-Q*  40.59  0.73 0.60 14.64  37.90  1.25 0.63 15.87  42.55  0.73 0.56 14.97
Ours  37.06  0.82 0.62 15.35  33.68  1.32 0.65 16.56  39.75  0.78 0.57 15.59
Table 2: Performance of LRQ-DiT (PixArt-$\alpha$) text-to-image generation on different datasets.
We follow DuQuant to construct an approximate rotation matrix $\hat{\mathbf{R}}$ for suppressing channel-wise outliers. First, the channel with the largest outlier concentration, $c^{*}$, is identified. A rotation matrix is greedily updated by swapping $c^{*}$ to the first channel and applying an orthogonal rotation $\mathbf{Q}$:
$\mathbf{R}^{i}=\mathbf{E}_{c^{*}}\,\tilde{\mathbf{R}}\,\mathbf{Q}\,\mathbf{E}_{c^{*}},\quad\hat{\mathbf{R}}=\prod_{i=1}^{n}\mathbf{R}^{i}$ (12)
where $\mathbf{E}_{c^{*}}$ is a switch matrix exchanging the first and $c^{*}$ columns, and $\tilde{\mathbf{R}}$ is an initialized orthogonal matrix. We perform $n$ greedy updates on the top-$k$ influential channels to obtain $\hat{\mathbf{R}}_{1}$, which mitigates outliers more effectively than a fixed Hadamard matrix. To further balance inter-block outlier distributions, channels are reordered in a staggered manner by sorting their numerical ranges and alternately assigning them across blocks, denoted as an orthogonal matrix $\mathbf{P}$. After permutation, another rotation $\hat{\mathbf{R}}_{2}$ is applied to further smooth residual outliers, effectively improving low-bit quantization performance. Our adaptive rotation scheme effectively mitigates both types of outliers while balancing computational cost and quantization performance. Notably, the entire strategy is training-free, avoiding expensive optimization for hundreds of linear layers in DiT models. Using the calibration data, we apply additional rotations to about 5% of layers that exhibit Salient Outliers. Simply setting the threshold to 1 yields satisfactory results in practice.
4 Experiment
4.1 Experimental Setting
Models, Datasets and Metrics. Following prior studies (Zhao et al. 2024; Li et al. 2024b), we evaluated performance across different DiT models: PixArt-$\Sigma$ (steps=20, CFG=4.5), PixArt-$\alpha$ (steps=20, CFG=4.5), FLUX.1-schnell (steps=4, CFG=0.), and FLUX.1-dev (steps=50, CFG=3.5). We selected 1024 prompts from COCO (Chen et al. 2015), MJHQ-30K (Li et al. 2024a), and the summarized Densely Captioned Images (sDCI) (Urbanek et al. 2024) respectively to evaluate the generalization capability of our method LRQ-DiT. We evaluated both the quality and similarity of images generated by the quantized model. Specifically, we utilize the Fréchet Inception Distance (FID, the lower the better) (Heusel et al. 2017) to measure the distributional distance between images generated by the quantized and FP16 models. Then, we employ Image Reward (IR, the higher the better) (Xu et al. 2023) to estimate human preference ratings for the generated images. We adopt Peak Signal Noise Ratio (PSNR, the higher the better) to measure the numerical similarity of the generated images. Structural Similarity Index Measure (SSIM, the higher the better) is used to measure the luminance, contrast, and structure similarity of the images generated by the quantized and FP16 models.
Precision Method FLUX.1-schnell 4 Steps FLUX.1-dev 50 Steps
COCO dataset MJHQ dataset sDCI dataset COCO dataset MJHQ dataset sDCI dataset
FID $\downarrow$ IR $\uparrow$ FID $\downarrow$ IR $\uparrow$ FID $\downarrow$ IR $\uparrow$ FID $\downarrow$ IR $\uparrow$ FID $\downarrow$ IR $\uparrow$ FID $\downarrow$ IR $\uparrow$
W3A4 Smooth* 71.66 0.48 78.35 0.36 72.35 0.19 156.64 -1.23 146.33 -1.15 166.03 -1.72
QuaRot* 60.46 0.85 65.09 0.84 63.73 0.68  98.37  0.14  96.68  0.31  96.78 -0.27
ViDiT-Q* 59.07 0.87 63.71 0.87 62.25 0.73  91.18  0.29  90.29  0.47  91.01 -0.07
Ours 50.77 0.92 54.52 0.98 56.40 0.79  58.81  0.57  63.13  0.76  63.58  0.25
W3A6 Smooth* 42.61 0.91 46.80 1.13 45.32 0.92  41.74  0.85  44.87  1.15  45.71  0.81
QuaRot* 41.52 0.99 44.97 1.24 44.40 1.03  39.73  0.94  43.21  1.22  44.14  0.94
Ours 41.47 0.99 40.81 1.26 43.74 1.03  38.60  0.94  39.57  1.22  43.31  0.95
W3A8 Smooth* 41.93 0.90 46.67 1.14 44.65 0.92  41.06  0.87  43.72  1.16  44.85  0.84
QuaRot* 40.74 1.00 44.62 1.24 43.68 1.01  39.20  0.94  41.97  1.22  43.94  0.93
Ours 41.36 0.98 40.43 1.25 43.56 1.04  38.48  0.95  39.12  1.23  42.83  0.93
Table 3: Performance of LRQ-DiT (FLUX.1) text-to-image generation on different datasets.
Methods COCO W3A4 COCO W3A6 COCO W3A8 MJHQ W3A4 MJHQ W3A6 MJHQ W3A8
TLQ ARS FID $\downarrow$ IR $\uparrow$ FID $\downarrow$ IR $\uparrow$ FID $\downarrow$ IR $\uparrow$ FID $\downarrow$ IR $\uparrow$ FID $\downarrow$ IR $\uparrow$ FID $\downarrow$ IR $\uparrow$
$\times$ $\times$ 245.37 -2.04 151.59 -1.46 149.14 -1.43 232.04 -1.86 124.88 -0.87 121.65 -0.82
$\times$ $\checkmark$ 237.15 -1.88 147.70 -1.19 140.53 -1.10 212.84 -1.65 120.88 -0.83 117.77 -0.83
$\checkmark$ $\times$  99.50 -0.29  65.61  0.00  64.87  0.01  76.02  0.38  55.14  0.61  54.85  0.62
$\checkmark$ $\checkmark$  90.36  0.01  57.76  0.31  56.95  0.30  75.51  0.48  51.95  0.85  50.84  0.87
Table 4: Ablation studies of TLQ and ARS under different quantization settings and datasets on PixArt-$\alpha$.
Implementation Details and Baselines.
We have aligned the experimental settings of ViDiT-Q (Zhao et al. 2024), which include employing per-channel static quantization for weights and per-token dynamic quantization for activations, along with the migration strength of $0.5$ to accomplish difficulty transfer from activations to weights. Besides, we retain full precision for certain sensitive layers with a small number of parameters. Our PTQ method utilizes 4-10 prompts to obtain the smoothing factor, metric $\mathbf{J}$, rotation matrix $\mathbf{R}$, and permutation matrix $\mathbf{P}$ during the calibration phase, without the need for retraining the model or fine-tuning parameters. Therefore, it does not introduce additional overhead during the inference phase. Compared to the performance gain brought by our method, the additional computational overhead is acceptable. We compare LRQ-DiT with methods SmoothQuant (Xiao et al. 2023), QuaRot (Ashkboos et al. 2024), and the current state-of-the-art PTQ method ViDiT-Q (Zhao et al. 2024) to validate the effectiveness of our method. To ensure fairness in the comparison of experimental results, we rerun these methods on both PixArt and FLUX models using A6000 and A100 GPUs, respectively.
4.2 Main Results
We present the evaluation results of our method and comparative methods in Tables 1, 2, and 3, as detailed below:
PixArt. We have conducted validation for the text-to-image generation tasks on the COCO, MJHQ, and sDCI datasets using both PixArt-$\Sigma$ and PixArt-$\alpha$, and evaluated their performance under different settings: W3A4, W3A6, W3A8, W4A4, W4A6, and W4A8. The experimental results in Table 1 and 2 demonstrate that text-to-image generation cannot be achieved solely through a naive smoothing scheme. QuaRot mitigates the impact of outliers in the model, enabling the generation of moderately satisfactory images at high-bit width. However, once quantized to low-bit width, the model’s performance experiences a sharp decline. ViDiT-Q, as the current state-of-the-art PTQ method, performs well under 4-bit weight quantization, but the performance deteriorates once the weights are quantized to 3-bit. Our method achieves excellent performance levels under 3-bit weight quantization and further enhances the model’s performance under other quantization settings.
FLUX. Based on previous research findings (Li et al. 2024b) and our experimental observations, we have reached the following conclusion: due to FLUX inherent advantages, the models are capable of generating high-quality images even after low-bit quantization. Specifically, only using naive smoothing scheme to achieve 4-bit quantization, the model still exhibits great performance. We conducted 3-bit weight quantization experiments on the FLUX.1-schnell and FLUX.1-dev models and validated their performance on different datasets. The results in Table 3 demonstrate that our method enhances the model’s performance under the W3A4 quantization setting, with improvements observed in both the quality and similarity of the generated images.
Figure 3: Visual Comparison of Generated Images from PixArt-$\Sigma$ under W3A6 Quantization.
4.3 Ablation Study
We conducted ablation experiments in Table 4 to validate the effectiveness of TLQ and ARS. Without TLQ and ARS, the model exhibits poor performance, and the generated images are also lacking in information. When ARS is applied independently to address the impact of mild and salient outliers in the model, the model’s performance is enhanced. When TLQ is applied independently to achieve more precise quantization of weights, the model’s performance also improves significantly. When TLQ and ARS are applied together, the quantized model achieves optimal performance in both image quality and similarity, providing strong evidence of the synergistic effect between TLQ and ARS.
4.4 Visualization Comparison
We generated images using QuaRot, ViDiT-Q and ours respectively on same prompts, with the specific results shown in Figure 3. LRQ-DiT demonstrates significant advantages in text-to-image generation. The foreground details in the generated images are presented with clarity and specificity, the depiction of motion scenes is precise and vivid, and at the same time, character recognition becomes much easier.
5 Conclusion
This paper introduces an efficient and accurate post-training quantization method LRQ-DiT for DiT models. LRQ-DiT identifies two key issues in model quantization and proposes corresponding solutions to address them. Specifically, TLQ utilizes twin-log quantization to allocate more quantization intervals to dense regions in weights, thereby reducing quantization errors and achieving precise quantization. Then, ARS identifies mild and salient outliers in activations through the metric and suppresses the influence of different types of outliers using an adaptive rotation scheme. We conduct quantization experiments with different bit settings on PixArt and FLUX models, and then validate the performance of text-to-image generation tasks on the COCO, MJHQ, and sDCI datasets. The experimental results demonstrate that LRQ-DiT not only enables the model to achieve satisfactory performance under 3-bit weight quantization but also enhances the model’s performance under other quantization settings, thereby exhibiting significant advantages.
Appendix A Visual Comparison
[…]
We provide more visual comparison results, as follows:
  • Figure 4 shows PixArt-$\Sigma$ under W3A4 quantization.
  • Figure 5 shows PixArt-$\Sigma$ under W3A6 Quantization.
  • Figure 6 shows PixArt-$\Sigma$ under W3A8 Quantization.
  • Figure 7 shows PixArt-$\alpha$ under W3A4 Quantization.
  • Figure 8 shows PixArt-$\alpha$ under W3A6 Quantization.
  • Figure 9 shows PixArt-$\alpha$ under W3A8 Quantization.
Figure 4: Visual Comparison of Generated Images from PixArt-$\Sigma$ under W3A4 Quantization.
Figure 5: Visual Comparison of Generated Images from PixArt-$\Sigma$ under W3A6 Quantization.
Figure 6: Visual Comparison of Generated Images from PixArt-$\Sigma$ under W3A8 Quantization.
Figure 7: Visual Comparison of Generated Images from PixArt-$\alpha$ under W3A4 Quantization.
Figure 8: Visual Comparison of Generated Images from PixArt-$\alpha$ under W3A6 Quantization.
Figure 9: Visual Comparison of Generated Images from PixArt-$\alpha$ under W3A8 Quantization.